Campaigns in 2015 :

Short testbeam period in DESY (line T24) to test the Spiroc2d chip.

More information in AHCALSpiroc2dTestBeam2015


We had two consequtive testbeam periods at SPS (CERN). One in July (from the 7th to the 21th July) and other in August (from 12th to 26th). A report of all the 2015 tb campaign (with special focusing in the summer test beams) was presented by K. Krueger at the CALICE meeting held at Munich.

Setup and operation summary report for the first period, see K. Krueger's slides and J. Kvasnicka's pictures of the installation.

Setup and operation summary report for the second period (?)

List of runs can be found here.

Please, go to the AHCALTestBeamCERN2015 wiki page where you can find more exhaustive and complete information about the daily operation, the data taking and storage, the data quality monitoring and the offline analysis.

AHCALTestBeamCERN2015

  1. AHCAL Testbeam Campaign at CERN 2015 (July/August)

  2. Setup

    1. July 2015

    2. August 2015

    3. DAQ Hierarchy

    4. Trigger, T0, Cherenkov channels and spill signal

    5. Common DAQ: EUDAQ

  3. Data

    1. Data access

    2. July

    3. August

  4. Operation

    1. Shifter tasks

    2. Beam operation

    3. Labview operation

    4. EUDAQ operation

  5. Monitoring

  6. Offline Data Analysis

  7. Run List

  8. Database Tags


A week of Testbeam has been done beginning of June as a "dress rehearsal" for CERN next summer 2015. This testbeam was operated in the EUDET Steel Stack, with a thickness of 17.2 mm.

The configuration was the following :

In front for the first Energy Scan Module with 2 Hamamatsu MPPCs with tranches, Module EBU_0, Module EBU_2, Module Mainz, Module 11, Module 10, Module 9, Module 8, Module 7, Module 6, Module 5, Module 12, Module 13, Module 14 and Module 15. The order is the same we had at Cern in July 2015.

Two Energy Scan was done from 1 GeV up to 5 GeV (with also collimator smaller to see beam energy spread) (The first energy scan contains good data but the Validation was not done right (late of 1 micro-sec)). A MIP Scan of EBU_0 was also done with the new settings provided by Katsu.

This Testbeam was a real success concerning data taking and also hardware (no stuck TDCs anymore). The quasi-online monitor could be checked and should be ready for the next period at CERN without major bugs. One new thing is that now a light tight box is on the top of the stack allowing us much more flexibility, also improvement on the cooling are done.

The new steering file and CalibrateProcessor will be available soon

The run list is available here.

Careful in using the data:

- there might be some corrupted bunch crossing ID.

- the new ITEP board (Module 11 & Module 10) might be in calib mode (Not in Physics mode) during the electrons run (it has been checked only on one run)


Another week of Testbeam at DESY. The setup was composed of 5 Layers in the Tungsten frames, everything in the Airstack. The Layer disposition was the following : Module 12, Module 13, Module 14, Modules ITEP (10 and 11 in a single slab in the top part of the frame!) and Module Mainz (also in the top part of the frame in the right corner).

The goal was to calibrate further the boards (specially ITEP and Mainz with new voltage settings), another goal was to understand further the stuck TDCs especially in the Module 14 and test the very first version of the quasi-online monitor. In overall, this was a success especially running with the EUDAQ, except for stuck TDCs. Concerning calibration, stability with the gain is now achieved and the Ketek Module are now at the right settings.

As usual, a list of remarks is available on the online elog here.


We had a TB period of about 1 week in February at DESY in a 3 GeV electron beam. The Setup was 5 big layers (Layer 12, 13, 14, 15 and one special layer composed of the 2 new ITEP boards and the SMD Mainz board) in an airstack configuration using the tungsten frames. The goal of this testbeam was to get a MIP Calibration for the outer tiles of theses layers and cross-check few things from CERN PS Testbeam.

In overall, the testbeam was a success. Thanks to it, we can get a light yield measure for the new Mainz board, cross-check results with the ones at CERN and get a better understanding of the electronics. One change was operated on the Layer 13. As we observed that the gain for that layer was too low at CERN, we decided to change the pre-amplifier values to lower values.

CERN PS

DESY Feb 2015

325fF

250fF

350fF

275fF -> 250fF

275fF

225fF

275fF

225fF

Daily runs of LED and Pedestal were made to check stability.

Remark : The reconstruction software from October/November CERN TB still works just the database paths and the processor Ahc2Calibrate needs to be modified!!! The source file is available here : Ahc2CalibrateProcessor

A list of remarks, comments and runs is available here.

airstack.jpg


CategoryHCAL

tb2015 (last edited 2017-01-05 13:04:46 by AmbraProvenza)